2 particles are projected upwards with speed 20m/s and 100m/s from the top of a tower. If the time of flight is in the ratio 1:3,then find the ht. of tower.

Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
relatively
-h=20T1-5T1^2=>5T1^2-20T1-h=0
-h=100T2-5T2^2=>5T2^2-100T2-h=0
=>T1/T2=(20+root(400+20h))/(100+root(10000+20h))=1/3
Now please evaluate yourself for h
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty